# No-hair and almost-no-hair results for static axisymmetric black holes and ultracompact objects in astrophysical environments

### v1

@@ Abstract
No-hair theorems are uniqueness results constraining the form of the metric of black holes in general relativity. These theorems are typically formulated under idealized assumptions, involving a mixture of local (regularity of the horizon) and global aspects (vacuum spacetime and asymptotic flatness). This limits their applicability to astrophysical scenarios of interest such as binary black holes and accreting systems, as well as their extension to horizonless objects. A previous result due to Gürlebeck constrains the asymptotic multipolar structure of static spacetimes containing black holes surrounded by matter although not revealing the possible structure of the metric itself. In this work, we disentangle some of these assumptions in the static and axisymmetric case. Specifically: i) we show that only a one-parameter family of black-hole geometries is compatible with a given external gravitational field, ii) we also analyze the case in which the central object is close to forming an event horizon but is still horizonless and show that the deviations from the natural black-hole shape have to die off as one approaches the black hole limit under the physical principle that curvatures are bounded.
@@ I Introduction
On the one hand, black holes are rarely isolated; they typically exist in (quasi-)equilibrium states surrounded by some form of matter or external gravitational fields. Although it is widely believed that such black holes remain “hairless”, it is not always crystal clear what is understood by that, namely, concerning the question of whether only a two-parameter family of black holes would be compatible with the structure of GR. To the best of our knowledge, the only results reported in that direction are those of Gürlebeck Gürlebeck (2014, 2015). Gürlebeck showed, in the static and axisymmetric case, that the Weyl multipole moments that characterize a spacetime close to infinity can be expressed as a sum of the contributions from black hole regions and regions with a non-vanishing energy-momentum tensor. The contributions from black holes are those associated with the Schwarzschild black holes. That is, a distorted black hole does not modify the contribution to the multipolar structure of the external gravitational field in which it is embedded. However, for instance, there is not a statement of a no-hair theorem asserting the existence of a single family of distorted black hole solutions given an external gravitational background.
On the other hand, the original formulation of the no-hair theorem requires the presence of an event horizon, specifically an infinite-redshift surface. However, recent results Barceló et al. (2019); Raposo et al. (2019) show that demanding the curvature at a surface of arbitrarily high redshift to be bounded and imposing vacuum Einstein equations up to spatial infinity places strong bounds on the multipolar structure of the spacetime. Furthermore, the multipole structure of the configuration approaches that of a black hole as the redshift of the surface parametrically goes to infinity. Hence, these results indicate that the requirement of an infinite-redshift surface can be relaxed without altering the overall physical picture.
@@
Here is an outline of the article. In Section II we review Israel’s theorem and revisit the different assumptions made in its derivation, with special emphasis on disentangling local and global aspects of the theorem. In Section III we present the main results of the article. In Subsec. III.1 we discuss the different vacuum solutions that exhibit an infinite redshift at some region and argue that only one of them leads to a smooth event horizon. In Subsecs. III.2 and III.3 we illustrate the difference between deforming a horizon from inside (some source or modification of GR inside the high-redshift surface) and outside (external gravitational fields). In Section IV we state a no-hair result for static and axisymmetric black holes in external gravitational fields based on the former analysis. In Section V we apply this machinery to the set-up in which there exists a maximum redshift surface with bounded curvatures and find bounds to the deviation of the multipole structure with respect to the case in which the redshift goes to infinity (the black hole limit). Finally, we conclude in Section VI with a summary of our work and comment on some aspects of no-hair theorems in astrophysical contexts and the interplay between our results and the vanishing of Love numbers. In Apps. AB and C, we have collected some technical derivations and statements of some known results that have been skipped in the main text for clarity of the presentation.
@@
In this article, we use the signature (,+,+,+)(-,+,+,+)( - , + , + , + ) for the spacetime metric and we work in geometrized units in which $c=G_{N}=1$. For the curvature tensors we use the conventions in the book of Wald (1984) i.e., $[\nabla_{A},\nabla_{B}]V^{C}=:-R_{ABD}{}^{C}V^{D}$, $R_{AB}:=R_{ACB}{}^{C}$. We use capital latin indices from the beginning of the alphabet as indices in the 4-dimensional manifold $(A,B...)$, lower case indices from the middle of the alphabet $(i,j...)$ for indices within the 3-dimensional submanifolds defined by constant time slices, and lower case indices from the beginning of the alphabet $(a,b...)$ for indices in 2-dimensional surfaces within these 3-dimensional submanifolds.
@@ II Revisiting Israel’s theorem: local vs global aspects
  • Take $\Sigma$ to be any $t=\text{constant}$ spatial hypersurface that is regular, empty, noncompact and asymptotically flat, namely that there exists a set of coordinates (that are asymptotically identified with the coordinates $x^{i}$) in which:
@@
  • The equipotential surfaces $\mathscr{V}=\text{constant}>0,$ $t=\text{constant}$ are regular, simply connected, and closed 2-spaces.
@@
Actually, let us show that the vacuum Einstein equations are compatible with any geometry induced at the horizon, as long as it is topologically a 2-sphere (times an open interval representing time). We can show this through a constructive procedure following the approach in Frolov and Sanchez (1986) that we sketch in here. We conveniently choose the coordinates $(t,\mathscr{V},\vec{\theta})$ so that the metric can be written as follows:
@@
in such a way that the horizon is located at $\mathscr{V}\rightarrow 0^{+}$. Then, it is possible to write down the Einstein vacuum equations in a neighbourhood of $\mathscr{V}=0$ for $\kappa$ and $h_{ab}$ and solve them for a series expansion of the form:
@@
The initial data required to solve these equations is only $\kappa^{(0)}$, which actually needs to be a constant (in agreement with the zeroth law of black hole dynamics Bardeen et al. (1973)) and corresponds to a choice of units, and $h_{ab}^{(0)}(\vec{\theta})$ which represents the induced metric on the horizon as a two-surface. Locally, there always exists a solution for every regular tensor $h_{ab}^{(0)}(\vec{\theta})$ and hence it is possible to have a local black hole with any kind of horizon shape. One can continue integrating outwards starting from one of these arbitrarily shaped horizons. However, Israel’s theorem tells us that only for the purely spherically symmetric shape, integrating outwards using Einstein vacuum equations leads to an asymptotically flat spacetime. We will show this explicitly in the next section, in which we restrict our attention to the case of static and axisymmetric solutions, where it is possible to provide sharper and clearer results.
@@ III Static and axisymmetric geometries
This form can always be achieved by means of coordinate transformations since the Lie algebra spanned by these two vectors is trivial, $[\bm{k},\bm{m}]=0$ Carter (1970).
@@
In Newtonian gravity, given some sources, one only needs to solve Poisson equation on a flat 3D space to obtain the gravitational potential $U$. Any source is compatible with the equation, for example, two masses at rest at a fixed distance. Poisson equation by itself does not tell you whether a configuration of sources that interact only gravitationally will be consistent or not with staticity. However, general relativity is a constrained system. One cannot give arbitrary conditions but only those compatible with the constraints. In this specific case this general issue appears in the solutions of the equation for $V$. As we will see, even if $U$ is perfectly regular it might happen that the associated $V$ exhibits irregularities. These irregularities inform us about some inconsistency that the selected source might have. In particular, in the absence of matter, the function $V$ must vanish on the $z$-axis, i.e. $V(0,z)=0$, to avoid the presence of conical singularities. This ensures that the length of a proper orbit of the axial Killing vector $\bm{m}$ that passes through a point $q$ located at a proper distance $d$ from the axis (fixed points of $\bm{m}$) is $2\pi d$ at leading order (see Mars and Senovilla (1993) for a rigorous demonstration).
@@
[…]
@@
[…]
@@
[…]
@@ III.1 Putative black hole spacetimes
For static metrics, an event horizon is a Killing horizon Hawking and Ellis (2023) and hence, at a horizon, the redshift function $e^{U}$ should vanish. This means that the function $U$ for which we need to solve an ALP, needs to go to $\infty$- ∞ at any putative horizon. Thus the horizon would not be covered by these coordinates. The regions at which the function $U$ goes to $\infty$- ∞ must correspond to those in which there is some kind of singular ($\delta$-like distributional) source. However, we highlight that a proper horizon itself would be in vacuum, and these sources would become an artifact of the ALP.
@@
As the ALP is a linear equation, we can always separate the effects of the different putative black holes present in the system. Thus, let us discuss first the elementary cases in which there is a just one of these singular behaviours, namely a point-like distribution at the axis, a ring-like distribution around the symmetry axis, or a constant density rod-like distribution sitting on the axis. One could think of a fourth kind of distribution, corresponding to performing an identification of the coordinates in the symmetry axis lying at the ends of the rod, leading to a toroidal black hole. However, these black holes require matter that violates the energy conditions in the outside region, and we will not specifically focus on them (see Appendix A, where we have included a description of these black holes for the sake of completeness).
@@
where $\mathcal{R}^{4}(R,\theta;M)$ is a function constructed from $M,R$ and $\sin\theta$, with dimensions of length to the fourth power:
@@
Clearly it displays a directional behaviour that is different from the Schwarzschild solution. For any value of $\theta\not\in\{0,\pi\}$ the Kretschmann scalar diverges as we approach $R\rightarrow 0$ since it has the behaviour
@@
where $N$ is a positive number and we are omitting the subdominant term in the argument of the exponential. For $\theta\in\{0,\pi\}$ the Kretchmann has a different behaviour and actually vanishes as we approach $R\rightarrow 0$ since in that case it exhibits a behaviour
@@
In both cases, the exponential behaviour dominates either to make the Kretschman scalar blow up or vanish. The equipotential surfaces and the surfaces of constant Kretschmann scalar are depicted in Fig. 3 (together with their Schwarzschild counterpart; see below).
Figure 3: We depict here in the upper panel the constant redshift surface in the $(r,z)$ coordinate for the Schwarzschild solution (left) and the Curzon solution (right), while we depict the Kretschmann scalar in the lower panel again for the Schwarzschild solution in the left and the Curzon solution to the right. We normalize the graphics so that the maximum is of ($\order{1}$ ).
@@
The Curzon metric exhibits another relevant feature. The area of the equipotential surfaces (constant $R$ surfaces) goes to infinity as $R\to 0$, see Fig. 4 for a representation of the area of constant $t$ and $R$ surfaces as a function of the radius $R$. Its explicit analytic expression can be found in Appendix C although it is not specially illuminating. This automatically excludes them from Israel’s theorem scope, due to the last assumption. Furthermore, it also shows that the putative horizon at $R\rightarrow 0$ is not regular in a different way. In fact, the directional behavior of the Kretschmann scalar, combined with the infinite extent of the constant $R$ region, suggests that the object, which appears point-like in these coordinates, actually represents an extended configuration.
@@
In fact, the behaviour of the geodesics in the Curzon metric was studied in Morgan and Szekeres (1973); Scott and Szekeres (1986a, b). Null geodesics approach the coordinate singularity $R=0$ in such a way that they enter parallel to the $z$-axis. Given that the surfaces $R=\text{constant}$ have an area that increases as we approach $R=0$, the surface $R=0$ is not really a point, but a whole plane of infinite extension. The geometry can be regarded as the end state of a very non-spherical anisotropic collapse which tends to compress all the matter into an infinitely diluted plane such that the matter content is moved away to infinity where a ring singularity is formed. Given that the plane is infinitely diluted, the curvature that is found by entering through the disk (which is what the null geodesics do in general, since they tend to enter towards the origin through the $z$-axis) is vanishing in an infinitely differentiable way, namely the $e^{-M/R}$ behavior (the archetypal example of a $C^{\infty}$ function which is not analytic). We can glue a flat spacetime to it which is what would be found by the light rays crossing the surface $R=0$ (which we emphasize is a plane, not a point). If we are to regard this as the end state of a collapse scenario, this is the only extension that we would do (as the continuation to the interior in black hole situations). If we consider an eternal configuration, we would also do an extension backwards in time (continue the outgoing geodesics to the past) and glue another flat spacetime region. Further continuations are allowed involving the behaviour of the other geodesics. Notice also that the extension is not unique, since the metric is not analytic, and hence there exists more (actually infinite) $C^{\infty}$ extensions of the metric across $R=0$.
We can also consider the behaviour of other singular pure multipole sources which are similar to the pure monopole behavior. The functions $U$ and $V$ in the metric for an $\ell$-order multipole configuration are Stephani et al. (2003)
$U^{(\ell)}=-\frac{M^{(\ell)}}{R^{\ell+1}}P_{\ell}\left(\cos\theta\right),$ (36)
@@
The Kretschmann scalar displays a behaviour which is qualitatively similar to the monopole one as we approach $R\rightarrow 0$, although with a more convoluted directional structure. Instead of having only two directions along which the Kretschmann vanishes, there are now a set of directions $\theta\in\{\theta_{i}\}_{i\in\mathcal{J}_{\ell}}$, with $\mathcal{J}_{\ell}$ a finite set for every $\ell$ along which the Kretschmann goes to zero. However, it is not easy to determine them in full generality and for every value of $\ell$. For any $\theta\neq\theta_{i}$, the leading behaviour of the Kretschmann is
@@
Consider the solution of the ALP to be that associated with an infinitely thin ring of matter located at $z=0$ and $r=r_{\rm ring}$, i.e. its distributional density is of the form $\rho(\bm{x})=M_{\text{ring}}\delta(z)\delta(r-r_{\rm ring})/2\pi$. This solution is often called the Bach-Weyl ring and is reviewed in detail in Appendix B. Axisymmetry implies that the ring has constant linear density along its circumference. There are several pathological features of the geometry associated with this solution.
First of all, close to the ring, the equipotential surfaces are toroids with smaller and smaller radii. This implies that they do not satisfy the hypotheses of Israel’s theorem. It is easy to see that the area of these equipotential toroids goes to zero as they approach the source since the vector generating axisymmetry $\bm{m}$ degenerates: The norm of $\bm{m}$ is given by $re^{U}$, and $r$ remains bounded whereas $e^{U}$ vanishes as $r\rightarrow r_{\rm ring}$. Thus, the area of the toroids also vanishes since it is proportional to $2\pi$ times the norm of $\bm{m}$ times the integral in the additional direction. This by itself tells us that they cannot approach a proper black hole horizon (of finite well-defined area) as it was noticed by Geroch and Hartle Geroch and Hartle (1982), which led them to exclude this possibility as potential black hole configurations in the static-axisymmetric setup.
@@
An additional pathology of these geometries that we have analyzed is the behaviour of the Kretschmann scalar. The Kretschmanm scalar diverges at the ring itself as it is illustrated in Fig. 5, leading to a singular geometry (which actually is of the naked-singularity type). Thus, even though they are excluded from Israel’s analysis by the assumption on the topology of constant redshift surfaces, we see that they cannot represent black holes with regular event horizons. This behaviour is generic of any singularity that is not located at the $z$-axis. Thus, we now focus on singularities located exclusively in the $z$-axis.
@@
The remaining simple type of source in the ALP to which we may associate an isolated putative black hole is that of a constant density rod located in the symmetry axis. Taking it to lie between $z=-M$ and $z=+M$, we have that the functions $U$ and $V$ for this source read:
@@
As is well known, this is precisely the Schwarzschild metric written in Weyl coordinates Stephani et al. (2003). The leading order behaviour of the functions $U$ and $V$ as $r\rightarrow 0$ for $z\in(-M,M)$ is $\log(r)$. Thus we automatically have the behaviour required for a putative black hole, namely,
@@
It has two crucial properties that might appear as surprising from the point of view of the Weyl coordinates, whereas they are well-known features of the Scwharzschild geometry. On the one hand, the area of the equipotential surfaces, which are ellipsoids in the fiduciary 3-space of the ALP, tends to a finite constant $A=16\pi M^{2}$ when these surfaces approach the rod. On the other hand, the Kretschmann scalar is constant in the equipotential surfaces and also goes to a finite value when the rod is approached. Therefore, the rod itself represents a regular geometrical boundary (i.e. a proper horizon) through which one can extend the geometry. We can see in this way that in this case the source of the ALP problem does not correspond to any kind of matter content from the general relativistic point of view: The Schwarzschild solution can be extended beyond the horizon as a vacuum solution Hawking and Ellis (2023). In the following sections we will analyze whether it is possible to deform the shape of this black hole by any means and in which cases do such perturbations lead to a singular behaviour.

### v3

@@ Abstract
No-hair theorems are uniqueness results constraining the form of the metric of black holes in general relativity. These theorems are typically formulated under idealized assumptions, involving a mixture of local (regularity of the horizon) and global aspects (everywhere vacuum spacetime and asymptotic flatness). This limits their applicability to astrophysical scenarios of interest such as binary black holes and accreting systems, as well as their extension to horizonless objects. A previous result due to Gürlebeck constrains the asymptotic multipolar structure of static spacetimes containing black holes surrounded by matter although not revealing the possible structure of the metric itself. In this work, we disentangle some of these assumptions in the static and axisymmetric case. Specifically: i) we show that only a one-parameter family of black-hole geometries is compatible with a given external gravitational field, ii) we also analyze the case in which the central object is close to forming an event horizon but is still horizonless and show that the deviations from the natural black-hole shape have to die off as one approaches the black hole limit under the physical principle that curvatures are bounded.
@@ I Introduction
On the one hand, black holes are rarely isolated; they typically exist in (quasi-)equilibrium states surrounded by some form of matter or external gravitational fields. Although it is widely believed that such black holes remain “hairless”, it is not always crystal clear what is understood by that, namely, concerning the question of whether only a two-parameter family of black holes would be compatible with the structure of GR. To the best of our knowledge, the only results reported in that direction are those of Gürlebeck Gürlebeck (2014, 2015). Gürlebeck showed, in the static and axisymmetric case, that the Weyl multipole moments that characterize a spacetime close to infinity can be expressed as a sum of the contributions from black hole regions and regions with a non-vanishing energy-momentum tensor. The contributions from black holes are those associated with the Schwarzschild black holes. That is, a distorted black hole does not modify the contribution to the multipolar structure of the external gravitational field in which it is embedded. However, for instance, there is no statement of a no-hair theorem asserting the existence of a single family of distorted black hole solutions given an external gravitational background.
On the other hand, the original formulation of the no-hair theorem requires the presence of an event horizon, specifically an infinite-redshift surface. However, recent results Barceló et al. (2019); Raposo et al. (2019) show that demanding the curvature at a surface of arbitrarily high redshift to be bounded and imposing vacuum Einstein equations up to spatial infinity places strong bounds on the multipolar structure of spacetime. Furthermore, the multipole structure of the configuration approaches that of a black hole as the redshift of the surface parametrically goes to infinity. Hence, these results indicate that the requirement of an infinite-redshift surface can be relaxed without altering the overall physical picture. All our arguments apply to the exterior region of black holes but not their interior region, which is also the case for the classic no-hair theorems.
@@
Here is an outline of the article. In Section II we review Israel’s theorem and revisit the different assumptions made in its derivation, with special emphasis on disentangling local and global aspects of the theorem. In Section III we present the main tools and results that will be used in the article. In Subsec. III.1 we discuss the different vacuum solutions that exhibit an infinite redshift at some region and argue that only one of them leads to a smooth event horizon. In Subsecs. III.2 and III.3 we illustrate the difference between deforming a horizon from inside (some source or modification of GR inside the high-redshift surface) and outside (external gravitational fields). In Section IV we state a no-hair result for static and axisymmetric black holes in external gravitational fields based on the former analysis. In Section V we apply this machinery to the set-up in which there exists a maximum redshift surface with bounded curvatures and find bounds to the deviation of the multipole structure with respect to the case in which the redshift goes to infinity (the black hole limit). Finally, we conclude in Section VI with a summary of our work and comment on some aspects of no-hair theorems in astrophysical contexts and the interplay between our results and the vanishing of Love numbers. In Apps. AB and C, we have collected some technical derivations and statements of some known results that have been skipped in the main text for clarity of the presentation.
@@
We use the signature (,+,+,+)(-,+,+,+) for the spacetime metric and we work in geometrized units in which $c=G_{N}=1$. For the curvature tensors we use the conventions in the book of Wald (1984) i.e., $\nabla_{A},\nabla_{B}]V^{C}=:-R_{ABD}{}^{C}V^{D}$, $R_{AB}:=R_{ACB}{}^{C}$. We use capital latin indices from the beginning of the alphabet as indices in the 4-dimensional manifold (A,B)(A,B...), lower case indices from the middle of the alphabet (i,j)(i,j...) for indices within the 3-dimensional submanifolds defined by constant time slices, and lower case indices from the beginning of the alphabet (a,b)(a,b...) for indices in 2-dimensional surfaces within these 3-dimensional submanifolds. The symbol $\Delta$ used at several places in the draft denotes a difference and not the Laplacian, which is always denoted by $\nabla^{2}$.
@@ II Revisiting Israel’s theorem: local vs global aspects
  • Take $\Sigma$ to be any $t=\text{constant}$ spatial hypersurface that is regular, empty, noncompact, and asymptotically flat, namely that there exists a set of coordinates (that are asymptotically identified with the coordinates $x^{i}$) in which:
@@
  • The equipotential surfaces $\mathscr{V}=\text{constant}>0,$ $t=\text{constant}$ are regular, simply connected, and closed 2-spheres.
@@
Actually, let us show that the vacuum Einstein equations are compatible with any geometry induced at the horizon, as long as it is topologically a 2-sphere (times an open interval representing time). We can show this through a constructive procedure following the approach in Frolov and Sanchez (1986) that we sketch in here. If an analytic static metric exists in an open neighborhood outside a regular topologically-spherical horizon (note that we are excluding more exotic mathematical situations such as the interesting nowhere-differentiable horizons discussed in Chrusciel and Galloway (1998)), then one can conveniently choose adapted coordinates $(t,\mathscr{V},\vec{\theta})$ so that the metric reads:
@@
with the horizon located at $\mathscr{V}\rightarrow 0^{+}$. Then, it is possible to write the Einstein vacuum equations in a neighborhood of $\mathscr{V}=0$ for $\kappa$ and $h_{ab}$ and solve them for a series expansion of the form:
@@
One can show that the consistency of this procedure requires $\kappa^{(0)}$ to be a constant (in agreement with the zeroth law of black hole dynamics Bardeen et al. (1973)), which corresponds to a choice of units. Apart from this condition, the Cauchy-Kovalevskaya theorem Evans (2010) ensures that there always exists a unique and analytical local solution for every analytic tensor $h_{ab}^{(0)}(\vec{\theta})$. Hence, it is possible to have a local black hole with any kind of horizon shape. Then, one can continue integrating outwards starting from one of these arbitrarily shaped horizons. However, Israel’s theorem tells us that only for the purely spherically symmetric shape, integrating outwards using the Einstein vacuum equations leads to an asymptotically flat spacetime. We will show this explicitly in the next section, in which we restrict our attention to the case of static and axisymmetric solutions, where it is possible to provide sharper and clearer results.
@@ III Static and axisymmetric geometries
This form can always be achieved by means of coordinate transformations since the Lie algebra spanned by these two vectors is trivial, $\bm{k},\bm{m}]=0$ Carter (1970).
@@
In Newtonian gravity, given some sources, one only needs to solve Poisson equation on a flat 3D space to obtain the gravitational potential $U$. Any source is compatible with the equation, for example, two masses at rest at a fixed distance. Poisson equation by itself does not tell you whether a configuration of sources that interact only gravitationally will be consistent or not with staticity. However, GR is a constrained system. One cannot give arbitrary conditions but only those compatible with the constraints. In this specific case this general issue appears in the solutions of the equation for $V$. As we will see, even if $U$ is perfectly regular it might happen that the associated $V$ exhibits irregularities. These irregularities inform us about some inconsistency that the selected source might have. In particular, in the absence of matter, the function $V$ must vanish on the $z$-axis, i.e. $V(0,z)=0$, to avoid the presence of conical singularities. This ensures that the length of a proper orbit of the axial Killing vector $\bm{m}$ that passes through a point $q$ located at a proper distance $d$ from the axis (fixed points of $\bm{m}$) is $2\pi d$ at leading order (see Mars and Senovilla (1993) for a rigorous demonstration).
@@
$V(r,z)=\int^{r}_{0}dr^{\prime}r^{\prime}\left[\left(\partial_{r^{\prime}}U(r^{\prime},z)\right)^{2}-\left(\partial_{z}U(r^{\prime},z)\right)^{2}\right].$ (20)
@@
$\int_{0}^{\infty}drr\left[\left(\partial_{r}U(r,z)\right)^{2}-\left(\partial_{z}U(r,z)\right)^{2}\right]=0.$ (21)
@@
$\oint_{\gamma}\bm{\omega}=\int_{S}dzdrr\partial_{z}U\nabla^{2}U=0.$ (25)
@@ III.1 Putative black hole spacetimes
For static metrics, an event horizon is a Killing horizon Hawking and Ellis (2023) and hence, at a horizon, the redshift function $e^{U}$ should vanish. This means that the function $U$ for which we need to solve an ALP, needs to go to -$\infty$ at any putative horizon. Thus the horizon would not be covered by these coordinates. The regions at which the function $U$ goes to -$\infty$ must correspond to those in which there is some kind of singular ($\delta$-like distributional) source. However, we highlight that a proper horizon itself would be in vacuum, and these sources would become an artifact of the ALP. Here, we focus exclusively on smooth horizons, specifically at least $\mathcal{C}^{2}$, to ensure the Einstein equations remain well-defined. However, beyond this framework, exotic horizons that are nowhere differentiable can be constructed Chrusciel and Galloway (1998), though we do not consider such cases in our analysis.
@@
As the ALP is a linear equation, we can always separate the effects of the different putative black holes present in the system. Thus, let us discuss first the elementary cases in which there is a just one of these singular behaviors, namely a point-like distribution at the axis, a ring-like distribution around the symmetry axis, or a constant density rod-like distribution sitting on the axis. One could think of a fourth kind of distribution, corresponding to performing an identification of the coordinates in the symmetry axis lying at the ends of the rod, leading to a toroidal black hole. However, these black holes require matter that violates the energy conditions in the outside region, and we will not specifically focus on them (see Appendix A, where we have included a description of these black holes for the sake of completeness).
@@
where $\mathcal{R}^{4}(R,\theta;M)$ is a function constructed from $M,R$ and $\theta$, with dimensions of length to the fourth power:
@@
Clearly it displays a directional behavior that is different from the Schwarzschild solution. For any value of $\theta\not\in\{0,\pi\}$ the Kretschmann scalar diverges as we approach $R\rightarrow 0$ since it has the behavior
@@
where $N$ is a positive number and we are omitting the subdominant term in the argument of the exponential. For $\theta\in\{0,\pi\}$ the Kretchmann has a different behavior and actually vanishes as we approach $R\rightarrow 0$ since in that case it exhibits a behavior
@@
In both cases, the exponential behavior dominates either to make the Kretschman scalar blow up or vanish. The equipotential surfaces and the surfaces of constant Kretschmann scalar are depicted in Fig. 3 (together with their Schwarzschild counterpart; see below).
[…]
@@
The Curzon metric exhibits another relevant feature. The area of the equipotential surfaces (constant $R$ surfaces) goes to infinity as $R\to 0$, see Fig. 4 for a representation of the area of constant $t$ and $R$ surfaces as a function of the radius $R$. Its explicit analytic expression can be found in Appendix C although it is not specially illuminating. This automatically excludes them from Israel’s theorem scope, due to the last assumption of the theorem. Furthermore, it also shows that the putative horizon at $R\rightarrow 0$ is not regular in a different way. In fact, the directional behavior of the Kretschmann scalar, combined with the infinite extent of the constant $R$ region, suggests that the object, which appears point-like in these coordinates, actually represents an extended configuration.
@@
In fact, the behavior of the geodesics in the Curzon metric was studied in Morgan and Szekeres (1973); Scott and Szekeres (1986a, b). Null geodesics approach the coordinate singularity $R=0$ in such a way that they enter parallel to the $z$-axis. Given that the surfaces $R=\text{constant}$ have an area that increases as we approach $R=0$, the surface $R=0$ is not really a point, but a whole plane of infinite extension. The geometry can be regarded as the end state of a very non-spherical anisotropic collapse which tends to compress all the matter into an infinitely diluted plane such that the matter content is moved away to infinity where a ring singularity is formed. Given that the plane is infinitely diluted, the curvature that is found by entering through the disk (which is what the null geodesics do in general, since they tend to enter towards the origin through the $z$-axis) is vanishing in an infinitely differentiable way, namely the $e^{-M/R}$ behavior (the archetypal example of a $C^{\infty}$ function which is not analytic). We can glue a flat spacetime to it which is what would be found by the light rays crossing the surface $R=0$ (which we emphasize is a plane, not a point). If we are to regard this as the end state of a collapse scenario, this is the only extension that we would do (as the continuation to the interior in black hole situations). If we consider an eternal configuration, we would also do an extension backwards in time (continue the outgoing geodesics to the past) and glue another flat spacetime region. Further continuations are allowed involving the behavior of the other geodesics. Notice also that the extension is not unique, since the metric is not analytic, and hence there exists more (actually infinite) $C^{\infty}$ extensions of the metric across $R=0$.
We can also consider the behavior of other singular pure multipole sources which are similar to the pure monopole behavior. The functions $U$ and $V$ in the metric for an $\ell$-order multipole configuration are Stephani et al. (2003)
$U^{(\ell)}=-\frac{M^{(\ell)}}{R^{\ell+1}}P_{\ell}\left(\cos\theta\right),$
@@
The Kretschmann scalar displays a behavior which is qualitatively similar to the monopole one as we approach $R\rightarrow 0$, although with a more convoluted directional structure. Instead of having only two directions along which the Kretschmann vanishes, there are now a set of directions $\theta\in\{\theta_{i}\}_{i\in\mathcal{J}_{\ell}}$, with $\mathcal{J}_{\ell}$ a finite set for every $\ell$ along which the Kretschmann goes to zero. However, it is not easy to determine them in full generality and for every value of $\ell$. For any $\theta\neq\theta_{i}$, the leading behavior of the Kretschmann is
@@
Consider the solution of the ALP associated with an infinitely thin ring of matter located at $z=0$ and $r=r_{\rm ring}$, i.e. a distributional density of the form $\rho(\bm{x})=M_{\text{ring}}\delta(z)\delta(r-r_{\rm ring})/2\pi$. This solution is often called the Bach-Weyl ring and is reviewed in detail in Appendix B. Axisymmetry implies that the ring has constant linear density along its circumference. There are several pathological features of the geometry associated with this solution.
First of all, as we approach the ring, the equipotential surfaces are toroids with smaller and smaller radii. This implies that they do not satisfy the hypotheses of Israel’s theorem. It is easy to see that the area of these equipotential toroids goes to zero as they approach the source since the vector generating axisymmetry $\bm{m}$ degenerates: The norm of $\bm{m}$ is given by $re^{U}$, and $r$ remains bounded whereas $e^{U}$ vanishes as $r\rightarrow r_{\rm ring}$. Thus, the area of the toroids also vanishes since it is proportional to $2\pi$ times the norm of $\bm{m}$ times the integral in the additional direction. This by itself tells us that they cannot approach a proper black hole horizon (of finite well-defined area) as it was noticed by Geroch and Hartle Geroch and Hartle (1982), which led them to exclude this possibility as potential black hole configurations in the static-axisymmetric setup.
@@
An additional pathology of these geometries that we have analyzed is the behavior of the Kretschmann scalar. The Kretschmanm scalar diverges at the ring itself as it is illustrated in Fig. 5, leading to a singular geometry (which is of the naked-singularity type). Thus, even though they are excluded from Israel’s analysis by the assumption on the topology of constant redshift surfaces, we see that they cannot represent black holes with regular event horizons. This behavior is generic of any singularity that is not located at the $z$-axis. Thus, we now focus on singularities located exclusively in the $z$-axis.
@@
The remaining type of source in the ALP to which we may associate an isolated putative black hole is that of a constant density rod located in the symmetry axis. Taking it to lie between z=Mz=-M and $z=+M$, we have that the functions $U$ and $V$ for this source read:
@@
As is well known, this is precisely the Schwarzschild metric written in Weyl coordinates Stephani et al. (2003). The leading order behavior of the functions $U$ and $V$ as $r\rightarrow 0$ for $z\in(-M,M)$ is $\log(r)$. Thus we automatically have the behavior required for a putative black hole, namely,
@@
It has two crucial properties that might appear as surprising from the point of view of the Weyl coordinates, whereas they are well-known features of the Scwharzschild geometry. On the one hand, the area of the equipotential surfaces, which are ellipsoids in the fiduciary 3-space of the ALP, tends to a finite constant $A=16\pi M^{2}$ when these surfaces approach the rod. On the other hand, the Kretschmann scalar is constant in the equipotential surfaces and also goes to a finite value when the rod is approached. Therefore, the rod itself represents a regular geometrical boundary (i.e. a proper horizon) through which one can extend the geometry. We can see in this way that in this case the source of the ALP problem does not correspond to any kind of matter content from the GR point of view: The Schwarzschild solution can be extended beyond the horizon as a vacuum solution Hawking and Ellis (2023). In the following sections we will analyze whether it is possible to deform the shape of this black hole by any means and in which cases do such perturbations lead to a singular behavior.